find the distance between the points (-20,16) and (-15,19). also find the midpoint of the segment that joins the points.
what is the distance?_____
what is the midpoint?______

Answer by meldocallos(2)

midpoint~ x=(-20 + -15)/2 = -35/2
y = (16 + 19)/ 2 = 35/2 so the midpoint is (-35/2, 35/2)
distance = square root of (-15 + 20)^2 + (19 - 16)^2 = sqrt of 34.